OpenStudy (zyberg):

What 3 digits should we add to number 523... (instead of dots), so that it would be divisible by 7, 8, 9?

OpenStudy (zyberg):

It really seemed to be a question for Chinese Remainder theorem, however, I don't think I can apply it here... 523000 + 100x + 10y + z. Using divisibility rules I could make a few equations: x + y + z = 8k 523000 + 100 x + 10y - 2z = 7n 10 + x + y + z = 9p Simplifying them a little bit I get: x + y + z = 8k 2 + 2x + 3y - 2z = 7n 1 + x + y + z = 9p However I have no idea what to do next...

OpenStudy (zyberg):

(oh, just n and p after simplification should be different numbers, sorry for my mistake!)

ganeshie8 (ganeshie8):

We should be able to do this simply by using divisibility rules

ganeshie8 (ganeshie8):

An integer is divisible by 9 if the sum of its digits are divisible by 9

OpenStudy (zyberg):

Well, yes, that's what I wrote with my equations: x + y + z = 8k //divisibility by 8 523000 + 100 x + 10y - 2z = 7n //divisibility by 7 10 + x + y + z = 9p //divisibility by 9

ganeshie8 (ganeshie8):

Ahh right, your equations represent the same..

ganeshie8 (ganeshie8):

stare at below two equations ``` x + y + z = 8k 1 + x + y + z = 9p ``` Is it easy to see that x+y+z has to equal 8 ?

ganeshie8 (ganeshie8):

Keep in mind that x, y, z are digits between 0 and 9

OpenStudy (zyberg):

Oh, I think that I had messed up with x + y + z = 8k... It should have been 100x + 10y + z = 8k.

ganeshie8 (ganeshie8):

Oh riht, I'm also not fully awake haha

OpenStudy (zyberg):

So, when simplifying it would get to 4x + 2y + z = 8n

OpenStudy (retireed):

Take 7*8*9 = 504 and find a number between 523000 and 523999 that is a multiple of 504 use excel or brute force

OpenStudy (zyberg):

@retirEEd the problem is that I am only allowed to use my head. ;)

OpenStudy (retireed):

That wasn't in the statement of the problem, so sorry I don't accept the response.

OpenStudy (jango_in_dtown):

Will the solution be always unique??

ganeshie8 (ganeshie8):

@Zyberg - I think the solution given by @retirEEd is the fastest way to solve this problem. No excel is needed however..

ganeshie8 (ganeshie8):

Start by finding 523000 mod 504

OpenStudy (zyberg):

352

OpenStudy (jango_in_dtown):

there will be 2 solutions then

ganeshie8 (ganeshie8):

That means 523000 + 352 is divisible by 504

ganeshie8 (ganeshie8):

Yeah looks two solutions

OpenStudy (zyberg):

ganeshie, shouldn't that be 523000 + 152?

OpenStudy (zyberg):

But dang, I see that it's very simple! So weird that I tried to solve it by using divisibility rules and such ;) Thank you as always, for pointing the right way to me! :)

ganeshie8 (ganeshie8):

Thank @retirEEd and please medal him he really helped us :)

ganeshie8 (ganeshie8):

Hey why are you adding 152 ?

OpenStudy (jango_in_dtown):

I fanned him... That method is the easiest

OpenStudy (zyberg):

Well, if 523000 is congruent to 352 in mod 504, that means that 523000 - 352 is divisible by 504. However, it is out of bounds, so we add 504 to it and we get -352+504 =152 ;)

OpenStudy (zyberg):

sure, after getting 523152 we would need to add 504 to the number to get the next one divisible by 504.

ganeshie8 (ganeshie8):

Wow I was so distracted. Thank you for explaining :)

OpenStudy (retireed):

The three numbers to add are 152 and I didn't use excel but it would have been easier then the using a simple calculator.